# Fast offline decoding with local message-passing automata

### v1

@@ 3 Message passing decoders
The automaton decoders studied in this work will be based on a particular type of message-passing scheme, where information about the spatial locations of anyons is transmitted using the classical variables $\mathbf{m}$. In the present subsection we describe the message-passing scheme for synchronous decoders operating on systems with periodic boundary conditions, which will be the setting considered in the majority of the main text. The Lindbladian extensions of these decoders are presented later in Sec. 6, and the extension to open boundary conditions is given in App. B.
We begin in Sec. 3.1 by explaining the construction in one dimension with periodic boundary conditions, where it performs error correction on the 1D repetition code. The generalization to higher dimensional systems with periodic boundary conditions is given subsequently in Sec. 3.2. Sec. 3.3 shows how our construction can be used to produce improved versions of the decoders in Ref. [13] (although these decoders do not perform better than our main construction). Finally, Sec. B.1 describes how to modify the decoder architecture to account for open boundary conditions.
@@ 3.1 Message passing in 1D
The feedback rules of the 1D decoder use the $m^{\pm}_{r}$ fields to attempt to move each anyon by one lattice site in the direction of the anyon nearest to it. Consider an anyon at site $r$. If $m^{+}_{r}=m^{-}_{r}=0$, the anyon has not received any messages, and no feedback is performed. If only $m^{\pm}_{r}$ is nonzero—so that a signal is received only from an anyon to the left (right) of the one at $r$—feedback is performed to move the anyon along the $\mp$ direction. Finally, if both $m^{\pm}_{r}\neq 0$, the anyon is moved along in the $\pm$ direction if $m^{\mp}_{r}<m^{\pm}_{r}$, since then the message coming from the $\pm$ direction was emitted earlier (hopefully, from a closer anyon) than the message coming from the $\mp$ direction. No anyon movement occurs if $m^{+}_{r}=m^{-}_{r}$.
@@ 3.2 Message passing in general dimensions
We now generalize the construction of the previous subsection to $D$ dimensions. The basic idea is the same as in one dimension, except now we introduce one message field for each of the $2D$ directions ±$\pm{\mathbf{\hat{a}}}$, $a=1,\dots,D$. We write $m_{\mathbf{r}}=\{m^{\pm a}_{r}\}$, where as before, each message field is valued in $\{0,1,\dots,m_{\sf max}\}$.
The automaton dynamics is defined so that $m^{\pm a}_{\mathbf{r}}$ propagates along the ±$\pm{\mathbf{\hat{a}}}$ direction, with the value of $m^{\pm a}_{\mathbf{r}}$ designed to convey the distance between $\mathbf{r}$ and the nearest anyon in the direction $-{\mathbf{\hat{a}}}$ from $\mathbf{r}$. In a bit more detail, the $m^{\pm a}$ messages emitted from an anyon at $\mathbf{r}$ are designed to travel outward and convey information about the anyon’s position to lattice sites in the cone-like region
@@
with the value of a message equal as before to the time over which that message has been propagating (see panel b of Fig. 3). $\mathcal{A}$ is designed such that the messages propagate over a lightfront dictated by the $\infty$ norm, with $m^{\pm a}_{\mathbf{r}}$ indicating the value of $||\mathbf{r}-\mathbf{r}^{\prime}||_{\infty}$, where $\mathbf{r}^{\prime}$ is the closest anyon to $\mathbf{r}$ in $\mathsf{C}^{\pm a}_{\mathbf{r}}$ (as in 1D, a value of $m^{\pm a}_{\mathbf{r}}=0$ indicates that no such anyon is present). These rules for $D=2$ are illustrated in panels b and c of Fig. 3, and are made precise in the following definition:
@@
denote the unit $\infty$-norm ball in the $D-1$ dimensions normal to ${\mathbf{\hat{a}}}$ centered at $\mathbf{r}$, and define $\mathsf{A}_{\Theta}(S)$ for a set $S\subset\Lambda$ as
@@
The automaton rule sources $m^{\pm a}_{\mathbf{r}}$ messages on the sites contained in $C^{(a)}_{\mathbf{r}\pm{\mathbf{\hat{a}}}}$, propagates them along the ±$\pm{\mathbf{\hat{a}}}$ direction, and increases their values so that $m^{\pm a}_{\mathbf{r}}$ is equal to the time at which the message at site $\mathbf{r}$ has been propagating, overwriting older messages with newer ones when appropriate. This is done using the rule
@@
$M_{\mathbf{r}}$ =$=\{(s,a)\in\{\pm 1\}\times\{1,\dots,D\}\,:\,m^{sa}_{\mathbf{r}}>0\},$ (26)
$(s_{\mathbf{r}},a_{\mathbf{r}})$ =$=\begin{dcases}{\sf argmin}^{\prime}_{(s,a)\in M_{\mathbf{r}}}\{m^{sa}_{\mathbf{r}}\}&M_{\mathbf{r}}\neq\varnothing\\ (0,0)&\text{else}\end{dcases},$
@@
${\sf FLF}(\mathbf{r},t)=\{(\mathbf{r}^{\prime},t)\,:\,t^{\prime}>t,\,v(t^{\prime}-t-1)<||\mathbf{r}-\mathbf{r}^{\prime}||_{\infty}\leq v(t^{\prime}-t)\}.$ (29)
The past lightfront is defined similarly as ${\sf PLF}(\mathbf{r},t)=\{(\mathbf{r}^{\prime},t^{\prime})\,:\,(\mathbf{r},t)\in{\sf FLF}(\mathbf{r}^{\prime},t^{\prime})\}.$ We will define the segment of the past lightfront along the ±$\pm{\mathbf{\hat{a}}}$ direction as
@@ 3.3 Other types of interactions
Suppose we want to consider a feedback rule that moves each anyon $a$ in a way that depends on the positions of the $N$ anyons closest to $a$. This can be done by using $N$ “flavors” of messages, with the control variables at each site being packaged as $m_{\mathbf{r}}=\{m^{\pm a,f}_{\mathbf{r}}\,:\,a\in\{1,\dots,D\},\,f\in\{1,\dots,N\}\}$. The update rules of $\mathcal{A}$ are then modified so that instead of newer messages overwriting older ones, a newer message of flavor $f$ turns an older message of flavor $f$ into a message of flavor $f+1$ (with a message of flavor $f=N$ simply being deleted). This allows the signals from up to $N$ anyons to simultaneously be stored at each site, at the cost of increasing the number of control variables at each site to ${\rm dim}(\mathcal{H}_{{\sf cl},\mathbf{r}})=\Theta(N\log\log(L))$. Since anyon pairs below threshold are very unlikely to be separated by a distance greater than $O(\log L)$, it suffices to set $N=\Theta((\log L)^{D})$ for any reasonable interaction scheme (taking $D=2$ then gives a slightly better scaling compared to the $\Theta((\log L)^{3}\log\log(L))$ required by the “constant-$c$” decoder of Ref. [13]).
@@ 4.2 Linear erosion
We should note that this property is not as strong as the linear erosion property discussed in the cellular automaton literature (see e.g. [34, 19, 35]), which would require erosion in both $\mathcal{H}_{\sf qu}$ and $\mathcal{H}_{\sf cl}$. In our decoders by contrast, erosion only occurs in $\mathcal{H}_{\sf qu}$. Indeed, the messages produced by anyons in a $(W,cW)$ cluster which propagate away from the cluster are not guaranteed to be erased by a time proportional to $W$, and thus have the potential of influencing anyons in an unboundedly large spacetime region. Nevertheless, the fact that the erosion time is linear in $\mathcal{H}_{\sf qu}$ will be shown to imply that these errant messages only slow down error correction by a constant factor (if erosion in $\mathcal{H}_{\sf qu}$ took longer than linear time, this effect would not be so innocuous, and would preclude the existence of a threshold). In passing, we note that the lack of erosion in $\mathcal{H}_{\sf cl}$ is ultimately the reason why these (and related) decoders fail to be robust online decoders; details will be provided in [27].
@@ 4.2.1 Modified message passing
To keep the proof of linear erosion simple, we will find it helpful to slightly modify the feedback rules applied by the decoder (proving linear erosion is the only place in this paper where these modified rules will be utilized). Let ${\mathbf{\hat{x}}}$ be the unit vector pointing along the first spatial coordinate $r^{1}$. Our modified feedback rules apply operators ~$\widetilde{\mathcal{F}}_{\mathbf{r}}^{(D)}$ which are obtained from the feedback operators $\mathcal{F}_{\mathbf{r}}^{(D)}$ of Def. 3.4 by adding motion of anyons along $-{\mathbf{\hat{x}}}$ to all feedback operators except those which move anyons along the ±$\pm{\mathbf{\hat{x}}}$ direction. Formally, this property is captured by the following definition:
@@
The modified message passing automaton decoder is defined as the decoder in Def. 3.4, except with feedback operators ~$\widetilde{\mathcal{F}}_{\mathbf{r}}^{(D)}$ defined as
~$\widetilde{\mathcal{F}}^{(D)}_{\mathbf{r}}$ $=X_{\langle\mathbf{r}-{\mathbf{\hat{x}}},\mathbf{r}\rangle}^{G_{\mathbf{r}}(\mathbf{m},{\bm{\sigma}})}\mathcal{F}^{(D)}_{\mathbf{r}},$ (38)
@@
where the $\widetilde{f}_{\mathbf{r}}^{a}$ are defined in the same way as in (26), except with a different way of resolving degeneracies: degeneracies between messages from the $s_{x}{\mathbf{\hat{x}}}$ and $s_{a}{\mathbf{\hat{a}}}$ directions with $a>1$ and $s_{x,a}\in\{\pm 1\}$ are split to favor motion along $s_{a}{\mathbf{\hat{a}}}$ if $a\neq 1$, while no anyon motion occurs if $s_{x}=+1$ (see panel a of Fig. 5 for an example in $D=2$). The function $G_{\mathbf{r}}(\mathbf{m},{\bm{\sigma}})$ is defined such that it equals 1 when both a) $\mathcal{F}^{(D)}$ moves an anyon along a link containing $\mathbf{r}$ and normal to ${\mathbf{\hat{x}}}$, and b) when an anyon remains at site $\mathbf{r}$ after the action of $\mathcal{F}^{(D)}$. Finally, note that this rule remains local since $\sigma_{\mathcal{F}^{(D)}(\mathbf{m},{\bm{\sigma}}),\mathbf{r}}$ only depends on $\{\mathbf{m},\mathbf{s}\}|_{B_{\mathbf{r}}^{(\infty)}(1)}$.
@@
Figure 5: a) modified feedback rules in 2D. The arrows in each of the four quadrants indicate the direction that an anyon moves upon receiving a signal from the associated direction. When the signals are received from the ±$\pm{\mathbf{\hat{y}}}$ components of the lightfront, the anyon moves along ±$\pm{\mathbf{\hat{y}}}$ before moving along $-{\mathbf{\hat{x}}}$. Degeneracies between different movement directions are resolved using the colors of the shaded lines on the boundaries of the lightfront segments; dashed black lines mean that anyons do not move. b) strategy of the erosion proof for clusters defined with the 1 norm. Anyons initially in a ball $\mathsf{B}$ of radius $W/2$ remain there if they are sufficiently isolated. Fixing the center of the ball at $\mathbf{r}={\mathbf{0}}$, after time $t_{\sf move}=W/v$, all anyons in $\mathsf{B}$ are confined to be to the left of the line $r^{1}=W/2-(t-t_{\sf move})$ (blue vertical lines); this leads to all anyons being annihilated after time $t_{\sf move}+W$.
@@ 4.2.2 Erosion of a single cluster: clean messages
This is true since we have assumed $v>1$: ${\sf PLF}^{+x}(\mathbf{r},t)\cap\mathsf{S}_{k}\neq\varnothing$ at time $t-(\delta-k)/v$ (again omitting floor functions), but our inductive assumption presumes that $\mathsf{S}_{k}$ was anyon-free starting at the earlier time $t-(\delta-k)<t-(\delta-k)/v$. Since $t>t_{\sf move}$, the anyon at $(\mathbf{r},t)$ must move. In particular, since $\mathsf{A}_{\Theta}({\sf PLF}^{+x}(\mathbf{r},t))=0$ and since our modified feedback rules always move anyons along $-{\mathbf{\hat{x}}}$ if the signal they receive does not come from ${\sf PLF}^{+x}$, this anyon must move from $\mathsf{S}_{\delta}$ to $\mathsf{S}_{\delta+1}$. Thus $\mathsf{S}_{\delta}$ must be anyon-free at time $t_{\sf move}+\delta+1$, which is what we wanted to show.
@@
$||\mathbf{r}_{b}-\mathbf{r}_{c}||_{\infty}\leq W<||\mathbf{r}_{a}-\mathbf{r}_{b}||_{\infty},$ (42)
@@
Consider first the case of a $(W,\infty)_{\infty}$-cluster. Let us first show that the anyons initially in $\mathsf{B}$ never leave $\mathsf{B}\cup\mathscr{P}^{x}_{\mathbf{r}-W{\mathbf{\hat{x}}}}(W/2)$. By a similar argument as in the proof of Lemma 4.4, one easily sees that due to the modified feedback rules, an anyon at $\mathbf{r}\in\partial\mathsf{B}$ can only move to $\Lambda\setminus\mathsf{B}$ if $\mathbf{r}$ is on the component of $\partial\mathsf{B}$ with unit normal along $-{\mathbf{\hat{x}}}$; the same argument shows that anyons at positions $\mathbf{r}\in\partial\mathscr{P}^{x}_{\mathbf{r}-W{\mathbf{\hat{x}}}}(W/2)\setminus\partial\mathsf{B}$ cannot move to $\Lambda\setminus\mathscr{P}^{x}_{\mathbf{r}-W{\mathbf{\hat{x}}}}(W/2)$. The time at which all anyons have received messages is upper bounded by $t_{\sf move}=\frac{3W}{2v}$ (the largest $\infty$-norm distance between two points in $\mathsf{B}\cup\mathscr{P}^{x}_{\mathbf{r}-W{\mathbf{\hat{x}}}}(W/2)$), which is a loose upper bound since the initial positions of the anyons are contained within $\mathsf{B}$. At $t\geq t_{\sf move}$, the same inductive argument as in the proof of Lemma 4.4 shows a direct analog of (40), viz.

### v4

@@ 3 Message passing decoders
The automaton decoders studied in this work will be based on a particular type of message-passing scheme, where information about the spatial locations of anyons is transmitted using the classical variables $\mathbf{m}$. In the present subsection we describe the message-passing scheme for synchronous surface code decoders operating on systems with periodic boundary conditions, which will be the setting considered in the majority of the main text. Extensions to general topological codes are given in sec. 5.4, the Lindbladian extensions of these decoders are presented in sec. 6, and open boundary conditions are treated in app. B.
We begin in sec. 3.1 by explaining the construction in one dimension with periodic boundary conditions, where it performs error correction on the 1D repetition code. The generalization to higher dimensional systems with periodic boundary conditions is given subsequently in sec. 3.2. Finally, sec. 3.3 shows how our construction can be used to produce improved versions of the decoders in Ref. [13].
@@ 3.1 Message passing in 1D
The feedback rules of the 1D decoder use the $m^{\pm}_{r}$ fields to attempt to move each anyon by one lattice site in the direction of the anyon nearest to it. Consider an anyon at site $r$. If $m^{+}_{r}=m^{-}_{r}=0$, the anyon has not received any messages, and no feedback is performed. If only $m^{\pm}_{r}$ is nonzero—so that a signal is received only from an anyon to the left (right) of the one at rr—feedback is performed to move the anyon along the $\mp$ direction. Finally, if both $m^{\pm}_{r}\neq 0$, the anyon is moved along in the $\pm$ direction if $m^{\mp}_{r}<m^{\pm}_{r}$, since then the message coming from the $\pm$ direction was emitted earlier (hopefully, from a closer anyon) than the message coming from the $\mp$ direction. No anyon movement occurs if $m^{+}_{r}=m^{-}_{r}$.
@@ 3.2 Message passing in general dimensions
We now generalize the construction of the previous subsection to $D$ dimensions. The basic idea is the same as in one dimension, except now we introduce one message field for each of the $2D$ directions $\pm{\mathbf{\hat{a}}}$, $a=1,\dots,D$. We write $m_{\mathbf{r}}=\{m^{\pm a}_{r}\}$, where as before, each message field is valued in $\{0,1,\dots,m_{\sf max}\}$.
The automaton dynamics is defined so that $m^{\pm a}_{\mathbf{r}}$ propagates along the $\pm{\mathbf{\hat{a}}}$ direction, with the value of $m^{\pm a}_{\mathbf{r}}$ designed to convey the distance between $\mathbf{r}$ and the nearest anyon in the direction $\mathbf{\hat{a}}$} from $\mathbf{r}$. In a bit more detail, the $m^{\pm a}$ messages emitted from an anyon at $\mathbf{r}$ are designed to travel outward and convey information about the anyon’s position to lattice sites in the cone-like region
@@
with the value of a message equal as before to the time over which that message has been propagating (see panel $\mathsf{b}$ of fig. 3). $\mathcal{A}$ is designed such that the messages propagate over a lightfront dictated by the $\infty$ norm, with $m^{\pm a}_{\mathbf{r}}$ indicating the value of $\mathbf{r}-\mathbf{r}^{\prime}||_{\infty}$, where $\mathbf{r}^{\prime}$ is the closest anyon to $\mathbf{r}$ in $\mathsf{C}^{\pm a}_{\mathbf{r}}$ (as in 1D, a value of $m^{\pm a}_{\mathbf{r}}=0$ indicates that no such anyon is present). These rules for $D=2$ are illustrated in panels b and c of fig. 3, and are made precise in the following definition:
@@
denote the unit $\infty$-norm ball in the D1D-1 dimensions normal to $^{\mathbf{\hat{a}}}$ centered at $\mathbf{r}$, and define $\mathsf{A}_{\Theta}(S)$ for a set $S\subset\Lambda$ as
@@
The automaton rule sources $m^{\pm a}_{\mathbf{r}}$ messages on the sites contained in $C^{(a)}_{\mathbf{r}\pm{\mathbf{\hat{a}}}}$, propagates them along the $\pm{\mathbf{\hat{a}}}$ direction, and increases their values so that $m^{\pm a}_{\mathbf{r}}$ is equal to the time at which the message at site $\mathbf{r}$ has been propagating, overwriting older messages with newer ones when appropriate. This is done using the rule
@@
$M_{\mathbf{r}}$ $=\{(s,a)\in\{\pm 1\}\times\{1,\dots,D\}\,:\,m^{sa}_{\mathbf{r}}>0\},$ (26)
[…]
@@
[…]
The past lightfront is defined similarly as ${\sf PLF}(\mathbf{r},t)=\{(\mathbf{r}^{\prime},t^{\prime})\,:\,(\mathbf{r},t)\in{\sf FLF}(\mathbf{r}^{\prime},t^{\prime})\}.$ We will define the segment of the past lightfront along the $\pm{\mathbf{\hat{a}}}$ direction as
@@ 3.3 Other types of interactions
Suppose we want to consider a feedback rule that moves each anyon $a$ in a way that depends on the positions of the $N$ anyons closest to $a$. This can be done by using $N$ “flavors” of messages, with the control variables at each site being packaged as $m_{\mathbf{r}}=\{m^{\pm a,f}_{\mathbf{r}}\,:\,a\in\{1,\dots,D\},\,f\in\{1,\dots,N\}\}$. The update rules of $\mathcal{A}$ are then modified so that instead of newer messages overwriting older ones, a newer message of flavor $f$ turns an older message of flavor $f$ into a message of flavor $f+1$ (with a message of flavor $f=N$ simply being deleted). This allows the signals from up to $N$ anyons to simultaneously be stored at each site, at the cost of increasing the number of control variables at each site to ${\rm dim}(\mathcal{H}_{{\sf cl},\mathbf{r}})=\Theta(N\log\log(L))$. Since anyon pairs below threshold are very unlikely to be separated by a distance greater than $O(\log L)$, it suffices to set $N=\Theta((\log L)^{D})$ for any reasonable interaction scheme (taking $D=2$ then gives a slightly better scaling compared to the $\Theta((\log L)^{3}\log\log(L))$ required by the “constant-cc” decoder of Ref. [13]).
@@ 4.2 Linear erosion
We should note that this property is not as strong as the linear erosion property discussed in the cellular automaton literature (see e.g. [35, 19, 36]), which would require erosion in both $\mathcal{H}_{\sf qu}$ and $\mathcal{H}_{\sf cl}$. In our decoders by contrast, erosion only occurs in $\mathcal{H}_{\sf qu}$. Indeed, the messages produced by anyons in a $(W,cW)$ cluster which propagate away from the cluster are not guaranteed to be erased by a time proportional to $W$, and thus have the potential of influencing anyons in an unboundedly large spacetime region. Nevertheless, the fact that the erosion time is linear in $\mathcal{H}_{\sf qu}$ will be shown to imply that these errant messages only slow down error correction by a constant factor (if erosion in $\mathcal{H}_{\sf qu}$ took longer than linear time, this effect would not be so innocuous, and would preclude the existence of a threshold). In passing, we note that the lack of erosion in $\mathcal{H}_{\sf cl}$ is ultimately the reason why these (and related) decoders fail to be robust online decoders; details are provided in Ref. [27].
@@ 4.2.1 Modified message passing
To keep the proof of linear erosion simple, we will find it helpful to slightly modify the feedback rules applied by the decoder (proving linear erosion is the only place in this paper where these modified rules will be utilized). Let $^{\mathbf{\hat{x}}}$ be the unit vector pointing along the first spatial coordinate $r^{1}$. Our modified feedback rules apply operators $\widetilde{\mathcal{F}}_{\mathbf{r}}^{(D)}$ which are obtained from the feedback operators $\mathcal{F}_{\mathbf{r}}^{(D)}$ of Def. 3.4 by adding motion of anyons along $\mathbf{\hat{x}}$} to all feedback operators except those which move anyons along the $\pm{\mathbf{\hat{x}}}$ direction. Formally, this property is captured by the following definition:
@@
The modified message passing automaton decoder is defined as the decoder in Def. 3.4, except with feedback operators $\widetilde{\mathcal{F}}_{\mathbf{r}}^{(D)}$ defined as
$\widetilde{\mathcal{F}}^{(D)}_{\mathbf{r}}$ $=X_{\langle\mathbf{r}-{\mathbf{\hat{x}}},\mathbf{r}\rangle}^{G_{\mathbf{r}}(\mathbf{m},{\bm{\sigma}})}\mathcal{F}^{(D)}_{\mathbf{r}}$, (38)
@@
where the $\widetilde{f}_{\mathbf{r}}^{a}$ are defined in the same way as in (26), except with a different way of resolving degeneracies: degeneracies between messages from the $s_{x}{\mathbf{\hat{x}}}$ and $s_{a}{\mathbf{\hat{a}}}$ directions with $a>1$ and $s_{x,a}\in\{\pm 1\}$ are split to favor motion along $s_{a}{\mathbf{\hat{a}}}$ if $a\neq 1$, while no anyon motion occurs if $s_{x}=+1$ (see panel a of fig. 5 for an example in $D=2$). The function $G_{\mathbf{r}}(\mathbf{m},{\bm{\sigma}})$ is defined such that it equals 1 when both a) $\mathcal{F}^{(D)}$ moves an anyon along a link containing $\mathbf{r}$ and normal to $^{\mathbf{\hat{x}}}$, and b) when an anyon remains at site $\mathbf{r}$ after the action of $\mathcal{F}^{(D)}$. Finally, note that this rule remains local since $\sigma_{\mathcal{F}^{(D)}(\mathbf{m},{\bm{\sigma}}),\mathbf{r}}$ only depends on $\{\mathbf{m},\mathbf{s}\}|_{B_{\mathbf{r}}^{(\infty)}(1)}$.
@@
Figure 5: a) modified feedback rules in 2D. The arrows in each of the four quadrants indicate the direction that an anyon moves upon receiving a signal from the associated direction. When the signals are received from the $\pm{\mathbf{\hat{y}}}$ components of the lightfront, the anyon moves along $\pm{\mathbf{\hat{y}}}$ before moving along $\mathbf{\hat{x}}$}. Degeneracies between different movement directions are resolved using the colors of the shaded lines on the boundaries of the lightfront segments; dashed black lines mean that anyons do not move. b) strategy of the erosion proof for clusters defined with the 1 norm. Anyons initially in a ball $\mathsf{B}$ of radius $W/2$ remain there if they are sufficiently isolated. Fixing the center of the ball at $\mathbf{r}={\mathbf{0}}$, after time $t_{\sf move}=W/v$, all anyons in $\mathsf{B}$ are confined to be to the left of the line $r^{1}=W/2-(t-t_{\sf move})$ (blue vertical lines); this leads to all anyons being annihilated after time $t_{\sf move}+W$.
@@ 4.2.2 Erosion of a single cluster: clean messages
This is true since we have assumed $v>1$: ${\sf PLF}^{+x}(\mathbf{r},t)\cap\mathsf{S}_{k}\neq\varnothing$ at time $t-(\delta-k)/v$ (again omitting floor functions), but our inductive assumption presumes that $\mathsf{S}_{k}$ was anyon-free starting at the earlier time $t-(\delta-k)<t-(\delta-k)/v$. Since $t>t_{\sf move}$, the anyon at $(\mathbf{r},t)$ must move. In particular, since $\mathsf{A}_{\Theta}({\sf PLF}^{+x}(\mathbf{r},t))=0$ and since our modified feedback rules always move anyons along $\mathbf{\hat{x}}$} if the signal they receive does not come from ${\sf PLF}^{+x}$, this anyon must move from $\mathsf{S}_{\delta}$ to $\mathsf{S}_{\delta+1}$. Thus $\mathsf{S}_{\delta}$ must be anyon-free at time $t_{\sf move}+\delta+1$, which is what we wanted to show.
@@
[…]
@@
Consider first the case of a $(W,\infty)_{\infty}$-cluster. Let us first show that the anyons initially in $\mathsf{B}$ never leave $\mathsf{B}\cup\mathscr{P}^{x}_{\mathbf{r}-W{\mathbf{\hat{x}}}}(W/2)$. By a similar argument as in the proof of Lemma 4.4, one easily sees that due to the modified feedback rules, an anyon at $\mathbf{r}\in\partial\mathsf{B}$ can only move to $\Lambda\setminus\mathsf{B}$ if $\mathbf{r}$ is on the component of $\partial\mathsf{B}$ with unit normal along $\mathbf{\hat{x}}$}; the same argument shows that anyons at positions $\mathbf{r}\in\partial\mathscr{P}^{x}_{\mathbf{r}-W{\mathbf{\hat{x}}}}(W/2)\setminus\partial\mathsf{B}$ cannot move to $\Lambda\setminus\mathscr{P}^{x}_{\mathbf{r}-W{\mathbf{\hat{x}}}}(W/2)$. The time at which all anyons have received messages is upper bounded by $t_{\sf move}=\frac{3W}{2v}$ (the largest $\infty$-norm distance between two points in $\mathsf{B}\cup\mathscr{P}^{x}_{\mathbf{r}-W{\mathbf{\hat{x}}}}(W/2)$), which is a loose upper bound since the initial positions of the anyons are contained within $\mathsf{B}$. At $t\geq t_{\sf move}$, the same inductive argument as in the proof of Lemma 4.4 shows a direct analog of (40), viz.